# RIPRAG: Hack a Black-box Retrieval-Augmented Generation Question-Answering System with Reinforcement Learning

### v1

@@ 5 Experiments and Analysis
  • RQ2: Can RIPRAG bypass state-of-the-art defense mechanisms designed for RAG systems?
@@ 5.1 Experiment Settings
To evaluate the generalizability of RIPRAG, we conduct experiments on three QA benchmarks: Natural Questions (NQ) [25], HotpotQA [26], and MS-MARCO [27]. The LLM we test on includes GLM4-9B [28], Qwen3-8B [18], InternLM2.5-7B-Chat [29], and DeepSeek-3.2-Exp [19]. The retriever we used in the target RAG system has two settings:
@@
    Naive retriever: Following the approach of PoisonedRAG, we adopt Contriever [30] as the retriever. In contrast, our approach retrieves ten times the number of candidate documents using Contriever and then reranks them with the BGE-M3 reranker [31], ultimately selecting the top-k documents as the retrieval results.
@@
    Complex retriever: In this setting, we construct a vector database using Milvus [32]. Specifically, we employ three retrieval components: the Qwen3-0.6B-Embedding model [33], the BGE-M3 embedding model, and the BGE-M3 reranker. We utilize the IVF_SQ8 index for efficient similarity search and apply Reciprocal Rank Fusion (RRF) to aggregate and refine the retrieval results into a final ranked list.
Since only PoisonedRAG and Topic-FlipRAG have proposed effective black-box attack methods so far, and Topic-FlipRAG differs from this study in terms of attacker capabilities and target definitions, we only adopt PoisonedRAG as the baseline. To demonstrate the limitations of white-box methods, we additionally include a fake white-box version of PoisonedRAG using Contriever as a replacement retriever for comparison.
@@ 5.2 Main Results (RQ1)
NQ
PoisonedRAG
(black-box)
0.48 0.35 0.29 0.52 0.32 0.32 0.60 0.46 0.39 0.39 0.26 0.23
PoisonedRAG
(fake white-box)
0.45 0.28 0.22 0.52 0.24 0.21 0.66 0.41 0.34 0.37 0.19 0.17
RIPRAG
(black-box)
0.70 0.72 0.94 0.72 0.89 0.76 0.85 89 0.62 0.42 0.38 0.52
HotpotQA
PoisonedRAG
(black-box)
0.71 0.54 0.53 0.75 0.51 0.55 0.82 0.60 0.59 0.65 0.43 0.39
PoisonedRAG
(fake white-box)
0.74 0.51 0.49 0.79 0.52 0.46 0.74 0.55 0.61 0.56 0.46 0.44
RIPRAG
(black-box)
0.88 0.87 1.00 0.82 0.97 0.94 0.95 0.93 0.86 0.70 0.56 0.55
MS-MARCO
PoisonedRAG
(black-box)
0.39 0.23 0.26 0.48 0.25 0.22 0.62 0.35 0.41 0.32 0.17 0.15
PoisonedRAG
(fake white-box)
0.36 0.20 0.18 0.38 0.15 0.17 0.52 0.28 0.22 0.26 0.17 0.12
RIPRAG
(black-box)
0.48 0.73 0.87 0.78 0.73 0.76 0.86 0.79 0.58 0.42 0.35 0.49
As shown in Table 1, RIPRAG significantly outperforms existing poisoning methods across diverse black-box RAG configurations. The proposed framework achieves substantially higher attack success rates (ASR) under both naive and complex retrieval settings, showcasing its robustness against varying system architectures and defense mechanisms. Notably, RIPRAG maintains strong performance even under challenging low-poisoning-rate scenarios where baseline methods exhibit severe degradation.
@@
RIPRAG establishes new state-of-the-art performance in black-box RAG poisoning, consistently surpassing both black-box and white-box variants of PoisonedRAG. RIPRAG achieves a maximum ASR improvement of 0.65 over PoisonedRAG (black-box) and 0.72 over PoisonedRAG (fake white-box) across all datasets and retrieval configurations. This substantial performance gap underscores the effectiveness of our RL-based optimization strategy, which systematically explores the black-box system’s preferences through iterative feedback rather than relying on static heuristics or surrogate models.
RIPRAG shows particular strength against complex RAG systems incorporating complex retrieval components, where gradient-based methods fail to work. Under the complex retrieval setting with hybrid searching, RIPRAG achieves 0.49-1.00 ASR across datasets. In contrast, PoisonedRAG’s performance deteriorates severely in these settings, with ASRs dropping to 0.12-0.59. This performance divergence highlights the limitation of methods that depend on simplified retrieval assumptions.
The framework demonstrates robust performance across different target RAG systems, indicating its general applicability beyond specific model architectures. On GLM4-9B, RIPRAG achieves 0.48-1.00 ASR across configurations, while maintaining 0.72-0.97 on Qwen3-8B, 0.58-0.95 on InternLM2.5-7B-Chat, and 0.35-0.70 on DeepSeek-v3.2-Exp. The consistent performance gains across all evaluated models confirm that RIPRAG’s effectiveness stems from its fundamental approach rather than model-specific optimizations.
RIPRAG exhibits exceptional effectiveness in low-poisoning-rate scenarios, overcoming a critical limitation of previous methods. When only a single poisoned document is injected (M=1), RIPRAG maintains high ASRs of 0.35-1.00 across datasets and models, while PoisonedRAG frequently shows performance collapse with ASRs as low as 0.12-0.61. This resilience stems from our RL agent’s ability to generate precisely optimized documents that maximize retrieval probability despite intense competition from legitimate documents. The performance preservation under constrained poisoning budgets demonstrates practical viability for real-world attack scenarios where massive document injection is infeasible.
Interestingly, RIPRAG sometimes achieves a higher ASR with M=1 than with M=3, suggesting the significant influence of batch size on policy optimization. With M=1, each query produces only one poisoned document, enabling a greater variety of queries per batch. This improves the accuracy of relative advantage estimation, leading to more stable policy gradient updates. In contrast, M=3 yields more documents per query but fewer distinct queries per batch, increasing noise in advantage estimation and potentially harming convergence.
Furthermore, the fake white-box version of PoisonedRAG performs worse than the black-box version, revealing its inherent limitations in real-world complex systems. For NQ with GLM4-9B under complex retrieval, fake white-box PoisonedRAG attains only 0.22 ASR versus 0.29 for black-box. The performance drop arises from two main causes: First, the fake white-box method depends on gradients from a simplified surrogate retriever, which misaligns with actual complex retrieval systems. Second, gradient-based text optimization methods like HotFlip often produce poisoned documents with grammatical or semantic flaws. Their poor readability and coherence reduce their likelihood of being trusted and used by the LLM, ultimately undermining attack success.
@@ 5.3 Defense Evaluation (RQ2)
As shown in Table 2, we conducted defense tests with the complex retriever setting and M=1. There are three defense methods: Rewriting query, HyDE [34], and RobustRAG [35]. RIPRAG maintains substantial ASR across all defense scenarios, ranging from 0.19 to 1.00. This consistent effectiveness underscores the adaptive capability of our RL-based approach, which learns to generate poisoned documents that remain effective even when defense mechanisms alter the retrieval or generation process.
@@
LLM GLM4-9B InternLM2.5-7B
NQ
N/A
0.94 0.62
Rewrite Query
0.76 0.51
HyDE
0.58 0.60
RobustRAG (M=3)
0.24 0.27
HotpotQA
N/A
1.00 0.86
Rewrite Query
0.94 0.78
HyDE
0.78 0.78
RobustRAG (M=3)
0.29 0.36
MS-MARCO
N/A
0.87 0.58
Rewrite Query
0.59 0.42
HyDE
0.41 0.33
RobustRAG (M=3)
0.19 0.14
The varying efficacy of different defense methods reveals critical insights about their operational mechanisms and limitations. Query Rewriting demonstrates moderate defensive capability, reducing RIPRAG’s ASR from 0.94 to 0.76 on NQ with GLM4-9B, but fails to provide comprehensive protection. This partial effectiveness stems from the method’s attempt to break potential semantic associations between queries and poisoned documents by reformulating user inputs. However, RIPRAG’s similarity reward component, which optimizes for broader semantic coherence rather than exact lexical matching, enables the generated documents to maintain relevance across multiple query formulations. The performance preservation indicates that our framework learns fundamental attack principles that transcend superficial textual variations.
RobustRAG emerges as the most effective defense, substantially reducing RIPRAG’s ASR to 0.19-0.29 across configurations, though complete mitigation remains elusive. Notably, due to the filtering strategy of RobustRAG, poisoning can never succeed when the number of poisoned documents is below the filtering threshold. Therefore, we used M=3 instead of M=1, where M is the same as the threshold in RobustRAG. However, RIPRAG can still achieve effective poisoning with 0.19-0.29 ASR even under extreme filtering conditions.
@@ 5.4 Ablation Study (RQ3)
Table 3: Contribution of different components in RIPRAG
NQ HotpotQA MS-MARCO
N/A
0.72 0.82 0.78
w. reference model
0.50 0.66 0.55
w.o. BRPO
0.17 0.61 0.14
w.o. similarity reward
0.09 0.24 0.06
w.o. attack reward
0.48 0.76 0.20
The comprehensive ablation study confirms that all components of RIPRAG contribute essentially to its overall effectiveness. As shown in Table 3, we conducted tests with the naive retriever setting and M=3, and the LLM is Qwen3-8B. The most dramatic drops occur when eliminating the similarity reward or BRPO.
@@
The similarity reward emerges as the most critical component for maintaining attack consistency, with its removal causing the most severe performance deterioration across all datasets. Without similarity guidance, ASR drops to 0.06 on MS-MARCO, 0.09 on NQ, and 0.24 on HotpotQA. The similarity reward serves as a dense training signal that creates a smooth optimization landscape that facilitates stable policy improvement throughout training.
@@ 5.5 Case Study
Through RLBF, RIPRAG not only mimics PoisonedRAG’s approach of embedding target questions into poisoned documents to increase retrieval probability but also actively acquires prompt injection techniques, effectively deceiving both retrieval and generation components. Here is an example of a poisoned document generated by RIPRAG:
@@ 6 Conclusion
In this work, we introduced RIPRAG, a novel black-box poisoning framework that leverages reinforcement learning to optimize adversarial documents against complex RAG systems. Our method significantly advances the state-of-the-art by demonstrating effective attacks without any internal knowledge of the target system, utilizing only binary success feedback to guide policy optimization. Through extensive experiments, we validated RIPRAG’s superiority over existing approaches across diverse datasets, model architectures, and retrieval configurations. The framework’s resilience against state-of-the-art defenses highlights critical vulnerabilities in current RAG security paradigms, while ablation studies confirmed the essential contributions of our technical innovations, particularly the BRPO algorithm and composite reward design.
@@ 7 Limitations
Despite its demonstrated effectiveness, RIPRAG possesses several limitations that warrant consideration. First, the framework requires substantial interaction with the target system during training, which may be impractical in scenarios with rate limitations or detection mechanisms. The RL optimization process typically necessitates hundreds to thousands of query-injection cycles to converge, potentially alerting system monitors to suspicious activity. Second, our approach assumes the attacker can successfully inject documents into the database, which may not be feasible in properly secured systems with rigorous content moderation. Finally, RIPRAG’s performance remains dependent on the quality and diversity of the initial query set, potentially limiting generalization to entirely unseen question types or domains not represented during training.
@@ Acknowledgement
This work is supported by the National Key R&D Program of China (2023YFC3306303), the Science and Technology Program of Zhejiang Province (2025C01087, 2023C01213), the Regional Innovation and Development Joint Fund of the National Natural Science Foundation of China (U22A6001), and Yongjiang Talent Introduction Program.

### v2

@@ 5 Experiments and Analysis
  • RQ2: To what extent can RIPRAG invalidate the defense measures of RAG systems?
@@ 5.1 Experiment Settings
To ensure a fair and rigorous evaluation of RIPRAG’s generalizability, our experiments are designed with a primary focus on equitable comparisons under controlled conditions. All methods are assessed on the same three widely-used QA benchmarks: Natural Questions (NQ) [21], HotpotQA [22], and MS-MARCO [23]. The target LLMs are held constant across all attacks and include GLM4-9B [24], Qwen3-8B [14], InternLM2.5-7B-Chat [25], and DeepSeek-3.2-Exp [15].
@@
    Naive retriever: We adopt Contriever [26] as the retriever following the approach of PoisonedRAG, and BGE-M3 reranker [27] as the reranker.
@@
    Complex retriever: To reflect production systems, we deploy a hybrid pipeline (Qwen3-0.6B-Embedding [28], BGE-M3 embedding/reranker) with Milvus [29] and RRF fusion.
We selected PoisonedRAG as the sole baseline because it is the only existing black-box method conforming to our threat model; all other state-of-the-art methods require white-box access to internal components of the target RAG system. For fair comparison, we exclusively benchmark against PoisonedRAG’s black-box configuration. Additionally, to expose the limitations of approaches that adapt white-box methods to black-box scenarios using open-source surrogate retrievers’ gradients, we include PoisonedRAG’s white-box variant as a reference and Contriever as the surrogate retriever, which we call the fake white-box variant of PoisonedRAG. All experiments maintain identical evaluation protocols to ensure comparability.
All experiments are evaluated under the same settings. In our experiments, the retrieval cut-off $k$ is set to 10, which is more challenging than the setting in PoisonedRAG. The consistent use of BM25 for similarity rewards simulates a realistic black-box scenario where attackers lack privileged access to the target system. This design choice guarantees that no method benefits from advantageous similarity modeling, thereby isolating the efficacy of the attack mechanisms themselves. While RIPRAG supports advanced neural rewards, we fix BM25 across all comparisons to maintain strict fairness.
@@ 5.2 Main Results (RQ1)
NQ PoisonedRAG (black-box) 0.48 0.35 0.29 0.52 0.32 0.32 0.60 0.46 0.39 0.39 0.26 0.23
PoisonedRAG (fake white-box) 0.45 0.28 0.22 0.52 0.24 0.21 0.66 0.41 0.34 0.37 0.19 0.17
RIPRAG (black-box) 0.70 0.72 0.94 0.72 0.89 0.76 0.85 0.89 0.62 0.42 0.38 0.52
HotpotQA PoisonedRAG (black-box) 0.71 0.54 0.53 0.75 0.51 0.55 0.82 0.60 0.59 0.65 0.43 0.39
PoisonedRAG (fake white-box) 0.74 0.51 0.49 0.79 0.52 0.46 0.74 0.55 0.61 0.56 0.46 0.44
RIPRAG (black-box) 0.88 0.87 1.00 0.82 0.97 0.94 0.95 0.93 0.86 0.70 0.56 0.55
MS-MARCO PoisonedRAG (black-box) 0.39 0.23 0.26 0.48 0.25 0.22 0.62 0.35 0.41 0.32 0.17 0.15
PoisonedRAG (fake white-box) 0.36 0.20 0.18 0.38 0.15 0.17 0.52 0.28 0.22 0.26 0.17 0.12
RIPRAG (black-box) 0.48 0.73 0.87 0.78 0.73 0.76 0.86 0.79 0.58 0.42 0.35 0.49
Table 1: Attack success rates (ASR) of different methods
@@
As shown in Table 1, RIPRAG significantly outperforms existing poisoning methods across diverse black-box RAG configurations, establishing a new state of the art. The framework achieves substantially higher attack success rates (ASR) under both naive and complex retrieval settings, with a maximum ASR improvement of 0.65 over PoisonedRAG (black-box) and 0.72 over PoisonedRAG (fake white-box). All results are averaged over 5 runs and exhibit highly stable performance, justifying the omission of variance in the table. This performance gap underscores the effectiveness of our RL-based optimization strategy, which systematically explores the black-box system’s preferences through iterative feedback rather than relying on static heuristics or surrogate models.
Notably, RIPRAG demonstrates particular strength against complex retrieval methods where gradient-based methods fail. Under hybrid searching, it maintains 0.49-1.00 ASR across datasets, while PoisonedRAG deteriorates to 0.12-0.59. The framework also exhibits robust generalization across different target LLMs, confirming that its effectiveness stems from a fundamental approach rather than model-specific optimizations.
A key advantage of RIPRAG is its resilience in low-poisoning-rate scenarios. With only a single poisoned document (M=1), it achieves 0.35-1.00 ASR, whereas PoisonedRAG frequently collapses to 0.12-0.61. This stems from the RL to generate precisely optimized documents that maximize poisoning ability. Interestingly, RIPRAG sometimes achieves a higher ASR with M=1 than with M=3, suggesting the significant influence of batch size on policy optimization. It might be because M=3 yields more documents per query but fewer distinct queries per batch, increasing noise in advantage estimation and potentially harming convergence. Furthermore, the fake white-box variant of PoisonedRAG underperforms even the black-box version, e.g., 0.22 vs. 0.29 ASR on NQ with GLM4-9B under complex retrieval. This reveals two inherent limitations: dependence on misaligned surrogate retriever gradients and the grammatical or semantic flaws introduced by gradient-based text optimization, which reduce document credibility and ultimately undermine attack success.
@@ 5.3 Defense Evaluation (RQ2)
As shown in Table 2, we conducted defense tests with the complex retriever setting and M=1, the target LLM is InternLM2.5-7B. There are three defense methods: Rewriting query, HyDE [30], and RAGuard [31]. RIPRAG maintains substantial ASR across all defense scenarios. This consistent effectiveness underscores the adaptive capability of our RL-based approach, which learns to generate poisoned documents that remain effective even when defense mechanisms alter the retrieval or generation process.
NQ HotpotQA MS-MARCO
Method PoisonedRAG RIPRAG PoisonedRAG RIPRAG PoisonedRAG RIPRAG
N/A 0.39 0.62 0.59 0.86 0.42 0.58
Rewrite Query 0.35 0.51 0.60 0.78 0.36 0.42
HyDE 0.32 0.60 0.58 0.78 0.33 0.33
RAGuard 0.06 0.10 0.11 0.13 0.11 0.16
RAGuard* 0.06 0.28 0.11 0.23 0.11 0.26
    Here RIPRAG is trained with doubled QLoRA rank.
@@
RAGuard emerges as the most effective defense, substantially reducing RIPRAG’s ASR to 0.10-0.16, though complete mitigation remains elusive. Notably, the limiting factor for RIPRAG’s ASR in this case is not the method itself, but rather the scale of trainable parameters. When we doubled the QLoRA rank, the ASR achieved nearly linear growth from 0.10-0.16 to 0.23-0.28.
@@ 5.4 Ablation Study (RQ3)
The comprehensive ablation study confirms that all components of RIPRAG contribute essentially to its overall effectiveness. As shown in Table 3, we conducted tests with the naive retriever setting and M=3, the target LLM is Qwen3-8B. The most dramatic drops occur when eliminating the similarity reward or BRPO.
Method NQ HotpotQA MS-MARCO
RIPRAG 0.72 0.82 0.78
w. reference model 0.50 0.66 0.55
w.o. BRPO 0.17 0.61 0.14
w.o. similarity reward 0.09 0.24 0.06
w.o. attack reward 0.48 0.76 0.20
Table 3: Contribution of components in RIPRAG
@@
The similarity reward emerges as the most critical component for maintaining attack consistency, with its removal causing the most severe performance deterioration across all datasets. Without similarity guidance, ASR drops to 0.06 on MS-MARCO, 0.09 on NQ, and 0.24 on HotpotQA. The similarity reward serves as a dense training signal that creates a smooth optimization landscape that facilitates stable policy improvement.
Figure 2: The impact of the retrieval cut-off $k$ on RIPRAG’s performance
We also analyzed the retrieval cut-off $k$. Figure 2 shows RIPRAG maintains robust ASR across all $k$, while PoisonedRAG degrades significantly. RIPRAG’s adaptive poisoning ensures effectiveness regardless of retrieval depth. In contrast, PoisonedRAG shows a sharp decline in ASR, indicating its vulnerability to increased retrieval depth. The divergence arises because PoisonedRAG relies on static poisoning that becomes diluted when more documents are retrieved, whereas RIPRAG dynamically optimizes poisoned content to preserve its prominence in the retrieval results. Notably, on MS-MARCO, RIPRAG’s ASR improves with larger $k$, suggesting its efficacy benefits from broader contextual coverage in certain domains. These results confirm that RIPRAG’s success is not constrained by the target RAG system’s retrieval depth, highlighting its practical applicability in real-world scenarios.
@@ 5.5 Case Study
Here is a successful poisoning case generated by RIPRAG.
@@
This case illustrates a sophisticated attack vector. The poisoned document begins with explicit instructions on how to manipulate the system to produce a specific wrong answer, effectively functioning as a meta-instruction for the RAG pipeline. It then presents factual birth dates but contradicts them with a false conclusion, claiming Lively was born earlier. This structure demonstrates that through RLBF, RIPRAG has evolved beyond merely embedding target questions into documents to boost retrieval likelihood. It now actively generates deceptive prompt-injection content that directly addresses and seeks to misguide the reasoning process of the system. Consequently, it poses a threat to both the retrieval stage (by increasing the document’s relevance score) and the generation stage (by injecting misleading logical instructions into the context), thereby compromising the entire RAG workflow.
6 Discussion
The proposed RIPRAG framework demonstrates superior attack performance across diverse RAG configurations, yet its reliance on RL raises concerns regarding computational costs. However, a practical analysis reveals that RIPRAG’s total cost is reasonable and often lower than gradient-based white-box methods. In our experiments, the fake white-box variant of PoisonedRAG consumed about 3 GPU hours, whereas RIPRAG required about 1 hour. Training RIPRAG against DeepSeek-V3.2-Exp incurred an API cost of about \$0.7 per instance. Although exceeding simple black-box heuristics, RIPRAG remains substantially cheaper than white-box alternatives, establishing it as a cost-effective solution for rigorous black-box security evaluation. Its one-time training yields a reusable policy for efficiently generating poisoned documents across new queries, justifying initial investment in real-world vulnerability assessment.
@@ 7 Conclusion
In this work, we introduced RIPRAG, a novel black-box poisoning framework that leverages reinforcement learning to optimize adversarial documents against complex RAG systems. Our method significantly advanced the state-of-the-art by demonstrating effective attacks without any internal knowledge of the target system, utilizing only binary success feedback to guide policy optimization. Through extensive experiments, we validated RIPRAG’s superiority over existing approaches across diverse datasets, model architectures, and retrieval configurations. The framework’s resilience against state-of-the-art defenses and low-poisoning-rate scenarios highlighted critical vulnerabilities in current RAG security paradigms. We also discussed the cost advantages and disadvantages of RIPRAG.
@@ 8 Limitations
Despite its demonstrated effectiveness, RIPRAG possesses several limitations that warrant consideration. First, the framework requires substantial interaction with the target system during training, which may be impractical in scenarios with rate limitations or detection mechanisms. Second, our approach assumes the attacker can successfully inject documents into the database, which may not be feasible in properly secured systems with rigorous content moderation. Finally, RIPRAG’s performance remains dependent on the quality and diversity of the initial query set, potentially limiting generalization to entirely unseen question types or domains not represented during training.
@@ Acknowledgement
This work is supported by the National Key R&D Program of China (2023YFC3306303), the National Natural Science Foundation of China (62502427), the Science and Technology Program of Zhejiang Province (2025C01087), the Yongjiang Talent Introduction Program (2024A-404-G), the Major Scientific and Technological Projects of CNTC(110202401031(SZ-05)), and the Zhejiang Key Laboratory Project (2024E10001).